Let f(x) = 2x + 1, then the number of real values of x for which the three unequal numbers f(x),f(2x), f(4x) are in geometric progression is

Grade:11

2 Answers

Sourabh Singh IIT Patna
8 years ago
Hii

Find out the values of the function and you will get the relation if you apply property of GM on the three given terms.
Divya Nitin Nikhade
4 years ago
(f(2x))power 2=(fx).(f(4x)) .                                            2(2x+1)power 2=(2x+1).(2(4x)+1) .                             (4x+1)power 2=(2x+1).(8x+1) .                                     16x power 2+1+8x=16xpower2+10x+1 .                     8x=10x .                                                                           10x-8x=0 .                                                                          2x=0 .                                                                               x=0